ADNAN MUHAMMED

Grade 12,

2. The length, breadth and height of a room are 5 m, 4 m and 3 m respectively. Find the cost of white washing the walls of the room and the ceiling at the rate of f 7.50 per m2

2. The length, breadth and height of a room are 5 m, 4 m and 3 m respectively. Find the cost of white washing the walls of the room and the ceiling at the rate of f 7.50 per m2

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Here, l = 5 m, b = 4 m, h = 3 m. Area of walls and ceiling = 2(1 + b) x h + lb = (2(5 + 4) x 3 + 5 x 4) m2 = (54 + 20) m2 = 74 m2 Cost = l' 7.50 x 74 = '556.